OpenStudy (emma.davenport):

which of the following equations is balanced? A. 4P+5O2 ------>2P2O5 B. P+O2 -------->P2O5 c. 4P+2O2 ------>2P4O5 D. P+2O2 ------->PO2

OpenStudy (emma.davenport):

@Abstract_Undercover @Englishguy @girlmeetsworld

OpenStudy (jfraser):

which equation has \(equal\) numbers of \(each \space type\) of atom on both sides of the arrow?

OpenStudy (whpalmer4):

I'll do one of them as an illustration of what JFraser said: \[\text{P}+\text{O}_2 \rightarrow\text{P}_2\text{O}_5\] On the left side we have: 1 P 2 O On the right side we have 2 P 5 O That's not balanced, as the numbers are not the same. A balanced version of that equation would be: \[4\text{P} + 5\text{O}_2\rightarrow 2\text{P}_2\text{O}_5\] Here we would have on the left side 4 P atoms and 5*2 = 10 O atoms On the right side we have 2 molecules each containing 2 P atoms and 5 O atoms, giving a total of 2*2 = 4 P atoms and 2*5 = 10 O atoms, which balances the left side exactly.